# GSRender: Deduplicated Occupancy Prediction via Weakly Supervised 3D Gaussian Splatting

### v1

@@ 3.5 Loss Function
$\mathcal{L}^{seg}_{adj}=-\sum_{i=1}^{n}\alpha_{\mathbf{s}(i)}\beta_{\mathbf{s}(i)}\mathbf{S}_{\text{pix}}(i)\log(\mathbf{\hat{S}}_{\text{pix}}(i)),$ (12)
$\mathcal{L}^{depth}_{adj}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\alpha_{\mathbf{s}(i)}d_{i}^{2}-\frac{1}{n^{2}}\left(\sum_{i=1}^{n}\alpha_{\mathbf{s}(i)}d_{i}\right)^{2}},$ (13)
where $d_{i}=\log\mathbf{D}_{i}-\log\hat{\mathbf{D}}_{i}$ denotes the difference between predicted and ground truth depth at pixel $i$, while $s(i)$ is the ground truth semantic category. $\alpha_{\mathbf{s}(i)}$ and $\beta_{\mathbf{s}(i)}$ represent the dynamic weight and category balance weight for $s(i)$, respectively. Consequently, our final loss function can be expressed as:
$\mathcal{L}=\mathcal{L}_{\text{curr}}+\sum_{k}\omega_{k}\mathcal{L}_{\text{adj},k}$ (14)
where $\omega$ represents the weight assigned to adjacent frames. Additionally, although we also considered using the distortion loss proposed in 2D Gaussian Splatting [13], we found it to be less effective and decided not to include it.
@@
BEVFormer[25] 3D 23.67 5.03 38.79 9.98 34.41 41.09 13.24 16.50 18.15 17.83 18.66 27.70 48.95 27.73 29.08 25.38 15.41 14.46
BEVStereo[21] 3D 24.51 5.73 38.41 7.88 38.70 41.20 17.56 17.33 14.69 10.31 16.84 29.62 54.08 28.92 32.68 26.54 18.74 17.49
OccFormer[53] 3D 21.93 5.94 30.29 12.32 34.40 39.17 14.44 16.45 17.22 9.27 13.90 26.36 50.99 30.96 34.66 22.73 6.76 6.97
SelfOcc[17] 2D* 9.30 - 0.15 0.66 5.46 12.54 0.00 0.80 2.10 0.00 0.00 8.25 55.49 - 26.30 26.54 14.22 5.60
OccNeRF[52] 2D* 9.54 - 0.83 0.82 5.13 12.49 3.50 0.23 3.10 1.84 0.52 3.90 52.62 - 20.81 24.75 18.45 13.19
GaussianOcc[10] 2D* 9.94 - 1.79 5.82 14.58 13.55 1.30 2.82 7.95 9.76 0.56 9.61 44.59 - 20.10 17.58 8.61 10.29
RenderOcc(1f)[35] 2D 19.33 1.82 19.54 12.6 16.08 12.85 8.76 12.14 9.33 12.94 17.60 12.13 56.25 28.83 36.89 38.82 15.08 16.98
GSRender(1f) 2D 21.36 2.44 20.13 13.56 19.54 18.38 8.35 18.77 10.67 14.81 16.63 14.92 61.74 29.77 35.95 40.56 19.83 17.06
RenderOcc[35] 2D+3D 26.11 4.84 31.72 10.72 27.67 26.45 13.87 18.20 17.67 17.84 21.19 23.25 63.20 36.42 46.21 44.26 19.58 20.72
GRender 2D+3D 29.56 8.42 34.93 15.12 30.71 29.61 16.70 9.48 17.61 16.58 23.92 27.24 77.94 39.21 51.69 54.61 23.82 24.92
Table 1: Quantitative Comparison on the Occ3D-nuScenes dataset in mIoU metric. 2D denotes the 2D labels generated using a semi-supervised method, rather than by projecting LiDAR point clouds onto the imaging plane. The best results obtained using only 2D supervision are highlighted in bold.
@@
FB-Occ (16f)[26] 3D R50 704×2absent2$\times$ 2× 256 90 33.5 26.7 34.1 39.7
SparseOcc(8f)[42] 3D R50 704$\times$256 24 34.0 28.0 34.7 39.4
Panoptic-FlashOcc (1f)[51] 3D R50 704$\times$256 24 35.2 29.4 36.0 40.1
SimpleOcc[9] 3D R101 672$\times$336 12 22.5 17.0 22.7 27.9
OccNeRF[52] 2D* R101 640$\times$384 24 10.5 6.9 10.3 14.3
GaussianOcc[10] 2D* R101 640$\times$384 12 11.9 8.7 11.9 15.0
RenderOcc(2f) 2D Swin-B 1408$\times$512 12 19.3 12.7 19.3 25.9
RenderOcc(7f) 2D Swin-B 1408$\times$512 12 19.5 13.4 19.6 25.5
GSRender(2f) 2D Swin-B 1408$\times$512 12 25.5(+6.0) 18.7(+5.3) 25.8(+6.2) 31.8(+6.3)
Table 2: Quantitative Comparison on the Occ3D-nuScenes dataset in RayIoU metric. The best results obtained using only 2D supervision are highlighted in bold.
@@
4 Experiments
Since the surround-view camera setup allows for better interaction between features from different perspectives and for a fair comparison with RenderOcc, we evaluated our method, GSRender, on the same datasets: NuScenes. Furthermore, we adopted a new evaluation metric, RayIoU[42], to highlight the effectiveness of our multi-frame compensation. Lastly, to gain deeper insights into GSRender, we conducted additional ablation studies and analysis on the NuScenes dataset.
4.1 Experimental Setup
Dataset. All of our experiments were conducted on the public dataset: NuScenes[3] and corresponding occupancy dataset, OCC3D-NuScenes[43], contains 600 scenes for training, 150 scenes for validation, covers a region centered around the ego-vehicle with dimensions [-40m, -40m, -1m, 40m, 40m, 5.4m] and a voxel size of [0.4m, 0.4m, 0.4m]. This dataset contains 18 classes, where classes 0 to 16 follow the same definitions as the nuScenes-lidarseg dataset, and class 17 represents free space. Since dataset does not provides 2D ground truth (GT), we followed the approach used in RenderOcc[35] to project the semantic and depth labels of 3D LiDAR points onto the 2D plane. This ensures a fair comparison with RenderOcc.
Implementation details. We implemented our model using PyTorch. We used off-the-shelf architecture Swin Transformer[29] as the image backbone with an input resolution of 512$\times$1408 and BEVStereo[21] for extracting 3D features. For the gaussian head, we simply employed a two-layer fully connected network. For the RC module, we used only one future frame as the basic experimental configuration for all experiments. During the training phase, we utilized the Adam optimizer with a linear learning rate scheduling strategy, set the batch size to 16, and trained for 12 epochs with a learning rate of 2e-4. All experiments were conducted on NVIDIA A6000 GPUs.
Evaluation metrics. SparseOcc [42] notes that overestimation can easily compromise the mIoU metric and introduces RayIoU to simulate LiDAR, enhancing scene completion evaluation through equal-distance resampling and temporal casting. Since the query ray measures distance based only on its first voxel contact, the model struggles to achieve a higher IoU with thick surfaces. Therefore, we utilized both the traditional mIoU metric and the newly proposed RayIoU metric to evaluate our method.
4.2 Performance
@@
We report the performance of GSRender in this part, demonstrating its superior results compared to the existing SOTA method, RenderOcc. We achieved significant performance improvements in RayIoU. Building on these enhancements, we now turn to a detailed comparison, both qualitative and quantitative, to further validate the efficacy of our approach.
4.2.1 Quantitative Comparison.
As illustrated in Table 1, We adhered to the original experimental setup of RenderOcc, conducting a comparative analysis between RenderOcc and GSRender under the condition of only using the current frame for supervision. The objective was to explore the performance of both methods when limited to a single-frame scenario, using mIoU as the evaluation metric. GSRender outperformed RenderOcc, achieving a higher mIoU (19.33 $\rightarrow$ 21.36), and enhanced performance when used as auxiliary supervision alongside 3D supervision, improving from 26.11 to 29.56.
@@
When supervised only by the current frame, the inherent limitations of 2D supervision cause both RenderOcc and GSRender to produce a thick and cracked surface. Consequently, relying solely on mIoU does not accurately reflect the model’s performance. To truly assess the capabilities of our model, we adopted RayIoU[42] as our primary evaluation metric to better evaluate the model’s true capability in scene reconstruction. As shown in Table 2, Notably, by utilizing only one frame for auxiliary supervision, our results showed a significant improvement over RenderOcc, which uses six auxiliary frames (19.5 $\rightarrow$ 25.5), even surpasses certain approaches that use 3D occupancy ground truth, such as SimpleOcc (22.5 $\rightarrow$ 25.5).
4.2.2 Qualitative Comparison.
@@
As illustrated in Figure 5, our experimental results demonstrate that the RC module significantly enhances the accuracy in positioning and shaping of specific categories, such as vehicles. In the third column, We visualized the results without using the RC module. This comparison clearly shows that reliance on single-frame ground truths tends to result in duplicated predictions. However, the integration of the RC module effectively addresses this problem, leading to a more unified and seamless reconstruction of the entire scene, enhanced the practicality of the prediction results.
4.2.3 Ablations and Analysis
@@
In our initial experiments, we postulated that using fixed Gaussian properties could also yield satisfactory rendering results. Nonetheless, this does not hold true. Therefore, we conducted ablation studies to verify the necessity of shifts in each Gaussian attribute, as shown in Table 3, indicate that Gaussian trained without(w/o) $\delta_{\mu},\delta_{s}$ and with(w) $\delta_{r}$ exhibit decreased accuracy. Furthermore, in Table 4, we also validated the effectiveness of the dynamic weight $\omega$ and the adjacent weight $\alpha$. Also presents the ablation experiment showing only the current frame as supervision, without the RC module. It can be seen that the RC module brings a significant improvement in performance. Notably, the mIoU for using only a single frame is comparable to that with the RC module, which is attributed to the hacked mIoU caused by duplicated predictions.
@@
Table 3: Ablation of Each Gaussian Properties
$RC$ $\omega$ $\alpha$ RayIoU(%)  mIoU(%)
@@
Table 4: Ablation of RC module and loss weights.
To explore the advantages of Gaussian representation, we filtered out empty Gaussians, then analyzed the distribution of mean shifts among the remaining Gaussians, as illustrated in Figure 6. Notably, the majority of these Gaussian means remained within their original positions, with shifts not exceeding half voxel. This stability enabled us during inference to directly associate the semantic labels of the Gaussians with their corresponding voxels, leading to robust results.
Furthermore, we visualized more intuitive outcomes, as demonstrated in the Figure 7, where the positions of Gaussian distributions were depicted as a point cloud. The second column displays results filtered by opacity alone. Many Gaussians in the sky lack supervision, which results in Gaussians near object surfaces being categorized as adjacent objects and exhibiting a tendency to shift closer to these objects. This also confirms the common z-offset in many Gaussians exceeds half a voxel and shows, as in Table 3, that overly constraining the Gaussian mean $\mu$ degrades performance. Building on this, as shown in the third column of Figure 7, we have additionally filtered out Gaussians that have shifted beyond the confines of their voxels. The results indicates that most Gaussians near real object surfaces adhered to their voxel positions. Additionally, Since the 2D ground truth is directly derived from LiDAR point clouds, it more accurately reflects the distribution of objects in the real world compared to 3D ground truth. Consequently, our method achieves object shapes , like trees, that are more realistic than the 3D ground truth.
@@
5 Limitation and future work
Qualitative results indicate that our method still has several limitations. We treat each voxel as a Gaussian distribution, but in practice, such a large number of redundant Gaussians is unnecessary, leading to increased computational burden and memory usage. Furthermore, the sparsity of LiDAR supervision could also impede the optimization of Gaussians, thereby affecting the overall performance of the model. GSRender serves as a foundation for our future work, where we will focus on reducing the number of Gaussians to further optimize performance.
6 Conclusion
@@
Supplementary Material
7 Additional Experiments and Analysis
@@
Table 5: Comparison of Different Frame Intervals. +1 represents the next immediate frame following the current frame.
@@
Table 6: Comparison of Different Frame Intervals. +1 represents the next immediate frame following the current frame
@@
Comparison of Different Frame Intervals. In the experimental section of the main text, we used only a single frame of future auxiliary supervision to simplify the experimental setup and focus on the core performance metrics. To gain a deeper understanding of how different auxiliary frame intervals impact the results, we conducted a series of additional experiments, as summarized in Table 6. These experiments systematically evaluated the performance using various combinations of auxiliary frames, including the next frame, the previous frame, the second previous frame, the second future frame, the previous two frames, the future two frames, and a combination of both previous and future two frames.
@@
8 Specific experimental configuration.
@@
9 Additional Limitations and Discussion
@@ 10 A Broader impacts
11 Licenses for involved assets
Our code is built on top of the codebase provided by RenderOcc[35] and BEVStereo[21], which is subject to the MIT license. Our experiments are carried out on the Occ3D-nuScenes[43] which provides occupancy labels for the nuScenes dataset[3]. Occ3D-nuScenes is licensed under the MIT license and nuScenes is licensed under the CC BY-NC-SA 4.0 license.

### v2

@@ IV-D Loss Function
$\mathcal{L}^{seg}_{adj}=-\sum_{i=1}^{n}\alpha_{s(i)}\beta_{s(i)}\mathbf{S}_{(i)}\log(\mathbf{\hat{S}}_{pix}(i)),$ (13)
$\mathcal{L}^{depth}_{adj}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\alpha_{s(i)}d_{i}^{2}-\frac{1}{n^{2}}\left(\sum_{i=1}^{n}\alpha_{s(i)}d_{i}\right)^{2}},$ (14)
where $d_{i}=\log\mathbf{D}_{i}-\log\hat{\mathbf{D}}_{i}$ denotes the difference between predicted and ground truth depth at pixel $i$, while $s(i)$ is the ground truth semantic category. $\alpha_{s(i)}$ and $\beta_{s(i)}$ represent the dynamic weight and category balance weight for $s(i)$, respectively. Consequently, our final loss function can be expressed as:
$\mathcal{L}=\mathcal{L}_{\text{curr}}+\sum_{k}\omega_{k}\mathcal{L}_{adj,k}$ (15)
where $\omega$ represents the weight assigned to adjacent frames. Additionally, although we also considered using the distortion loss proposed in 2D Gaussian Splatting [huang20242d], we found it to be less effective and decided not to include it.
V Experiments
We trained and evaluated our method on the nuscenes dataset, leveraging the surround-view camera setup for better feature interaction. Additionally, we introduced RayIoU [tang2024sparseocc] to highlight the effectiveness of multi-frame compensation and conducted ablation studies for deeper insights into GSRender.
@@ V Experiments
BEVFormer [li2022bevformer] 3D 23.67 5.03 38.79 9.98 34.41 41.09 13.24 16.50 18.15 17.83 18.66 27.70 48.95 27.73 29.08 25.38 15.41 14.46
BEVStereo [li2023bevstereo] 3D 24.51 5.73 38.41 7.88 38.70 41.20 17.56 17.33 14.69 10.31 16.84 29.62 54.08 28.92 32.68 26.54 18.74 17.49
OccFormer [zhang2023occformer] 3D 21.93 5.94 30.29 12.32 34.40 39.17 14.44 16.45 17.22 9.27 13.90 26.36 50.99 30.96 34.66 22.73 6.76 6.97
SelfOcc [huang2024selfocc] 2D* 9.30 - 0.15 0.66 5.46 12.54 0.00 0.80 2.10 0.00 0.00 8.25 55.49 - 26.30 26.54 14.22 5.60
OccNeRF [zhang2023occnerf] 2D* 9.54 - 0.83 0.82 5.13 12.49 3.50 0.23 3.10 1.84 0.52 3.90 52.62 - 20.81 24.75 18.45 13.19
GaussianOcc [gan2024gaussianocc] 2D* 9.94 - 1.79 5.82 14.58 13.55 1.30 2.82 7.95 9.76 0.56 9.61 44.59 - 20.10 17.58 8.61 10.29
RenderOcc(1f) [pan2024renderocc] LiDAR-2D 19.33 1.82 19.54 12.6 16.08 12.85 8.76 12.14 9.33 12.94 17.60 12.13 56.25 28.83 36.89 38.82 15.08 16.98
GSRender(1f) LiDAR-2D 21.36 2.44 20.13 13.56 19.54 18.38 8.35 18.77 10.67 14.81 16.63 14.92 61.74 29.77 35.95 40.56 19.83 17.06
RenderOcc [pan2024renderocc] LiDAR-2D+3D 26.11 4.84 31.72 10.72 27.67 26.45 13.87 18.20 17.67 17.84 21.19 23.25 63.20 36.42 46.21 44.26 19.58 20.72
GSRender LiDAR-2D+3D 29.56 8.42 34.93 15.12 30.71 29.61 16.70 9.48 17.61 16.58 23.92 27.24 77.94 39.21 51.69 54.61 23.82 24.92
TABLE I: Quantitative Comparison on the Occ3D-nuScenes dataset in mIoU metric. 2D denotes the semantic and depth maps generated using a semi-supervised method, rather than by projecting LiDAR point clouds onto the imaging plane. The best results obtained using only 2D supervision are highlighted in bold.
@@
[…]
TABLE II: Quantitative Comparison on the Occ3D-nuScenes dataset in RayIoU metric. The best results obtained using only LiDAR-2D supervision are highlighted in bold.
V-A Experimental Setup
V-A1 Dataset.
All of our experiments were conducted on the nuscenes dataset [caesar2020nuscenes] and the corresponding occupancy dataset, OCC3D-NuScenes [tian2024occ3d]. The occupancy dataset contains 600 scenes for training and 150 scenes for validation. It covers a region centered around the ego-vehicle with dimensions $40\text{m},-40\text{m},-1\text{m},40\text{m},40\text{m},5.4\text{m}]$ and uses a voxel size of $0.4\text{m},0.4\text{m},0.4\text{m}]$. This dataset contains 18 classes, where classes 0 to 16 follow the same definitions as the nuScenes-lidarseg dataset, and class 17 represents free space. We followed the approach used in RenderOcc [pan2024renderocc] to project the semantic and depth labels of 3D LiDAR points onto the 2D maps.
@@ V-A Experimental Setup
V-A2 Implementation details.
[…]
V-B Performance
@@
We report the performance of GSRender in this part, demonstrating its superior results compared to the existing SOTA method. We achieved significant performance improvements in RayIoU. Building on these enhancements, we now turn to a detailed comparison, both qualitative and quantitative, to further validate the efficacy of our approach.
V-B1 Quantitative Comparison.
@@
As illustrated in Table I, The objective was to explore the performance of both methods when limited to a single-frame scenario, using mIoU as the evaluation metric. GSRender outperformed RenderOcc, achieving a higher mIoU, and enhanced performance when used as auxiliary supervision alongside 3D supervision, improving from 26.11 to 29.56. When supervised only by the current frame, the inherent limitations of single view supervision cause both RenderOcc and GSRender to produce a thick surface. Consequently, relying solely on mIoU does not accurately reflect the model’s performance. To truly assess the capabilities of our model, we adopted RayIoU [tang2024sparseocc] as our primary evaluation metric to better evaluate the model’s true capability in scene reconstruction. As shown in Table 2. it is surprising that by utilizing only one frame for auxiliary supervision. Our results achieved a significant improvement over RenderOcc which uses six auxiliary frames. In fact our approach even surpasses certain methods that rely on 3D occupancy ground truth such as SimpleOcc.
V-B2 Qualitative Comparison.
@@
As illustrated in Figure 5, our experimental results demonstrate that the RC module significantly enhances the accuracy in positioning and shaping of specific categories, such as vehicles. In the third column, we visualized the results without using the RC module. This comparison clearly shows that reliance on single-frame ground truths tends to result in duplicated predictions. However, the integration of the RC module effectively addresses this problem, leading to a more unified and seamless reconstruction of the entire scene and enhancing the practicality of the prediction results.
V-B3 Ablations and Analysis.
@@
In our initial experiments, we postulated that using fixed Gaussian properties could also yield satisfactory rendering results. Nonetheless, this does not hold true. Therefore, we conducted ablation studies to verify the necessity of shifts in each Gaussian attribute, as shown in Table IV, indicate that Gaussian trained without mean and scale shift $\delta_{\mu},\delta_{s}$ and with rotation shift $\delta_{r}$ exhibit decreased accuracy. Furthermore, in Table IV, we also validated the effectiveness of the dynamic weight $\omega$ and the adjacent weight $\alpha$. Also presents the ablation experiment showing only the current frame as supervision, without the RC module. It can be seen that the RC module brings a significant improvement in performance. Notably, the mIoU for using only a single frame is comparable to that with the RC module, which is attributed to the hacked mIoU caused by duplicated predictions.
@@
TABLE III: Ablation of Each Gaussian Properties
RCRC $\omega$ $\alpha$ RayIoU(%)  mIoU(%)
@@
TABLE IV: Ablation of RC module and loss weights.
To better investigate the benefits of the Gaussian representation, we excluded Gaussians that do not correspond to any occupied space, then analyzed the distribution of mean shifts among the remaining Gaussians, as illustrated in Figure 6. Notably, the majority of these Gaussian means remained within their original positions, with shifts not exceeding half voxel. This stability enabled us during inference to directly associate the semantic labels of the Gaussians with their corresponding voxels, leading to robust results.
Furthermore, we visualized more intuitive outcomes, as demonstrated in the Figure 7, where the positions of Gaussian distributions were depicted as a point cloud. The second column displays results filtered by opacity alone. Many Gaussians in the sky lack supervision, which results in Gaussians near object surfaces being categorized as adjacent objects and exhibiting a tendency to shift closer to these objects. This also confirms the common z-offset in many Gaussians exceeds half a voxel and shows, as in Table III, that overly constraining the Gaussian mean $\mu$ degrades performance. Building on this, as shown in the third column of Figure 7, we have additionally filtered out Gaussians that have shifted beyond the confines of their voxels. The results indicates that most Gaussians near real object surfaces adhered to their voxel positions. Additionally, since the 2D ground truth is directly derived from LiDAR point clouds, it more accurately reflects the distribution of objects in the real world compared to 3D ground truth. Consequently, our method achieves object shapes , like trees, that are more realistic than the 3D ground truth.
@@
VI Limitation and future work
Qualitative results indicate that our method still has several limitations. We treat each voxel as a Gaussian, but in practice, such a large number of redundant Gaussians is unnecessary, leading to increased computational burden and memory usage. Furthermore, the sparsity of LiDAR supervision could also impede the optimization of Gaussians, thereby affecting the overall performance of the model. GSRender serves as a foundation for our future work, where we will focus on reducing the number of Gaussians to further optimize performance.
VII Conclusion
@@
VIII Additional Experiments and Analysis
@@
TABLE V: Comparison of Different Frame Intervals. +1 represents the next immediate frame following the current frame.
@@
TABLE VI: Comparison of Different Frame Intervals. +1 represents the next immediate frame following the current frame
@@
Comparison of Different Frame Intervals. In the experimental section of the main text, we used only a single frame of future auxiliary supervision to simplify the experimental setup and focus on the core performance metrics. To gain a deeper understanding of how different auxiliary frame intervals impact the results, we conducted a series of additional experiments, as summarized in Table VI. These experiments systematically evaluated the performance using various combinations of auxiliary frames, including the next frame, the previous frame, the second previous frame, the second future frame, the previous two frames, the future two frames, and a combination of both previous and future two frames.
@@
IX Specific experimental configuration.
@@
X Additional Limitations and Discussion
@@ XI A Broader impacts
XII Licenses for involved assets
Our code is built on top of the codebase provided by RenderOcc[pan2024renderocc] and BEVStereo[li2023bevstereo], which is subject to the MIT license. Our experiments are carried out on the Occ3D-nuScenes[tian2024occ3d] which provides occupancy labels for the nuScenes dataset[caesar2020nuscenes]. Occ3D-nuScenes is licensed under the MIT license and nuScenes is licensed under the CC BY-NC-SA 4.0 license.